OpenStudy (anonymous):

help ?!?!?!?!?!?!?!(:

OpenStudy (anonymous):

solve ?!?!?!?(: \[\frac{ x+1 }{ 2 }=\frac{ 7 }{ 5 }\]

OpenStudy (akashdeepdeb):

Cross multiply. Then solve like a simple equation.

OpenStudy (jonnyvonny):

When approached with 2 fractions that equal one another, you may "cross multiply". So, you would multiply the 2 and 7 and the (x+1) by 5. \[14=5x+5\] From then, you solve for x. \[14=5x+5\rightarrow9=5x \rightarrow9/5=x\] I hope this helped, feel free to pst me if you need further elaboration or assistance with any other problem. :)

OpenStudy (anonymous):

so \[x= \frac{ 11 }{ 3 }\] ???

OpenStudy (anonymous):

No. Try looking up what cross multiplying is and how to do it. This is what you do. (5)(x+1) = (7)(2) 5x + 5 = 14 5x = 14 - 5 5x = 9 x = 9/5 So the answer is x = 9/5.